### v3

# How glass breaks – Damage explains the difference between surface and fracture energies in amorphous silica

Abstract
The difference between free surface energy and fracture toughness in amorphous silica is studied via multi-scale simulations. We combine the homogenization of a molecular dynamics fracture model with a phase-field approach to track and quantify the various energy contributions. We clearly separate free surface energy localized as potential energy on the surface and damage diffusion over a 16-23 Å range around the crack path. The plastic contribution is negligible. These findings, which clarify brittle fracture mechanisms in amorphous silicates, align with toughness measurements in silica.
Surface free energy quantifies the excess energy associated with unbalanced interatomic bonds at a material’s surface. In solids, the bulk state is always energetically more favorable, so creating a surface requires energy input. Griffith [1] postulated a similar concept in his theory of crack propagation, stating that a crack can advance if the elastic energy released exceeds a material-specific threshold. Essentially, both free surface energy and fracture surface energy describe the same underlying phenomenon.
However, in experiments, fracture surface energy is almost always significantly higher. It is the case for polymers [2, 3] and crystalline metals [4] where the difference between free surface and fracture energies are found to span several orders of magnitude, because of plastic dissipation around the crack tip, and also for bulk metallic glasses [5, 6]. One of the rare exceptions is silicon crystals [7, 8, 9], where the difference falls within the precision limits of the methods used, and which can be assumed to break by individual bond rupture at an atomically sharp tip as expected from a brittle material [10],
Intriguingly, for silicate glasses, the archetypes of brittle materials, it has long been known that the measured fracture energy exceeds the surface free energy by a factor of ca. 5 [11, 12, 13]. The difference has been tentatively ascribed to plastic dissipation [14, 11] but as it is difficult to experimentally evidence plasticity at this scale in an amorphous material, let alone quantify it (and also because it openly conflicts with the accepted notion of brittleness), the claim has remained controversial [15] to these days: signs of plasticity at silica crack tips have been discounted [16] almost as soon as found [17]. Worse still, standard models for plastic dissipation in cracks [18] are inapplicable to silicates because their plasticity is essentially devoid of hardening [19].
Faced with this conundrum, we focused on silica glass to investigate the rupture of brittle amorphous silicates at the crossover between the atomic and the continuum scales. Previous atomic scale simulations have recognized the existence of an inelastic zone around the crack tip [20, 21]. Here, we show that we can clearly separate the two contributions to the fracture energy : 1) the energy required for surface formation—accounting for bond breaking and atomic relaxation near the surface—, and 2) the energy due to structural rearrangements extending much deeper into the material, over a distance of approximately 20 Å. We also show that the latter process is best described as damage rather than plasticity or nonlinear elasticity and that the ratio between the two contributions conforms to the experiments. This approach offers a fully consistent framework for understanding fracture in amorphous silica, an archetype of brittle materials.
Simulation methods.—To explore the fracture properties of amorphous silica, the atomistic simulations employed the BKS potential [22] to model atomic interactions and Wolf’s truncation method to handle Coulombic forces [23]. The BKS potential was shown to qualitatively but consistently describe the elastic and plastic mechanical properties of these systems [24, 25]. Amorphous glass samples were generated by randomly distributing atoms within a periodic box. Molecular dynamics simulations were then performed to equilibrate, quench, and test the samples. Athermal deformations, pertinent for systems below the glass transition temperature, were applied through iterative energy minimization using the conjugate gradient method. Further simulation details can be found in the Supplementary Materials.
Diffuse damage.—Fracture simulations were carried out on a 3D sample with dimensions $L_{x}\times L_{y}\times L_{z}=400\times 300\times 100$ Å 3, containing 800k atoms, which was found to be large enough to capture initiation and propagation adequately (Fig. 1). The initial defect, a rounded incision with a radius of $r_{c}=10$Å extends along the $z$ direction on a face normal to the $x$ direction. There is periodicity in the $z$ direction. For the boundary conditions on the $x$ and $y$ faces, a K-field displacement was imposed, simulating uniaxial tension along $y$ at infinity. The crack then initiated and propagated along the $x$ direction as the loading parameter ($K_{I}$) was increased. Further details of the fracture simulation are available in the Supplementary Material.
The atomic scale fields of the cracked sample were then coarse-grained (CG), which essentially involves Gaussian convolution to ensure mass and energy conservation during the discrete-continuum transition. The choice of the CG width $w$ will be discussed in detail below. As a result, the following continuum quantities were available in both Lagrangian and Eulerian configurations: displacements $u$, Hencky (logarithmic) strain tensor $\bm{\varepsilon}$, Cauchy stress tensor $\bm{\sigma}$ , potential energy density $\psi_{\rm{pot}}$ (sum of the atomic interactions) and mass density $\rho$.
Figure 1: Damage field obtained from molecular simulations: (a) Distribution of damage in the middle plane of the sample. (b) Circles represent the damage profile along the $y$ direction at $x=125$ Å, fitted with a Gaussian function of maximum height $d_{\rm{max}}$ and width $l$. (c) Variation of the fitted width $l$ along the crack and under different global loading states, color-coded by $d_{\rm{max}}$ according to the colorbar. The blue region indicates $l_{c}$ identified using the Finite Element Update (FEMU) scheme based on the phase-field formulation (see eq. 4).
Under the assumption of linear elasticity, we can calculate the elastic strain energy density field  [^1] Capital $\Psi$ is the global value integrated over the whole volume ($V$), while small $\psi$ refers to local energy densities. as
$\psi_{\epsilon}\leavevmode\nobreak\ =\leavevmode\nobreak\ {\psi_{\epsilon,0}}+\int_{{\bm{\varepsilon}_{0}}}^{\bm{\varepsilon}}{\bm{\sigma}:d\bm{\varepsilon}}\approx{\psi_{\epsilon,0}}+{\textstyle{1\over 2}}\bm{\sigma}:\bm{\varepsilon}J,$ (1)
where $J$ denotes the determinant of the deformation gradient tensor, accounting for volume changes, and $\psi_{\epsilon,0}$ is the initial strain energy density, which may be non zero, typically after quenching.
In the case of purely elastic deformation, the local elastic energy density field $\psi_{\epsilon}$ is equal to the undamaged strain energy $\psi_{0}$ calculated from the strain and the initial local elastic stiffness. In contrast, our results evidence a dramatic reduction of the local elastic energy density $\psi_{\epsilon}$ over a large area surrounding the crack faces. This result is evidence for a significant evolution of the structure of the material around the crack tip.
To quantify this evolution, we resort to the continuum mechanics concept of damage, where the local material stiffness decreases gradually down to a vanishing stiffness which signals fracture [27]. Damage can thus conveniently be described as a scalar field defined by:
${\psi_{\epsilon}}\left({{\bf{u}},d}\right)={\left({1-d}\right)^{2}}\psi_{0}\left({\bm{\varepsilon}\left({\bf{u}}\right)}\right).$ (2)
Thus $d$ evolves between 0 for the pristine and 1 for the fully damaged material. We have used the spectral energy decomposition scheme [28], i.e., the energy is calculated based on the principal values of the strain tensor, thereby preventing crack opening under compression. However, here, the compressive term is negligible and has been omitted from eq. (2). The detailed calculation can be found in the Supplementary Material.
Fig. 1a illustrates the damage distribution from molecular simulations in an $xy$ plane slice of the simulation box after fracture propagation over ca., 150 Å. Fig. 1b shows a cross section along $y$ at $x=125$ Å with a Gaussian fit from which the damage width $l$ is derived. Finally, Fig. 1c presents the damage width $l$ along the crack. The width varies but typically equals 20 Å along the crack faces. The color map indicates maximum damage values along the $y$ axis. Of course, for consistency, the CG width $w$ must be kept small compared to the width of the damage zone $l$. As shown in detail in the Supplementary Material, we found that for w=35w=3–5 Å, the mean value of $l$ remains the same, with only the fluctuation decreasing. At $w=8$ Å, the mean value starts to increase slightly, but this increase remains within the order of 10 %. As $w$ increases beyond 8 Å, the width of the damage zone, $l$, increases linearly with $w$. The value $w=8$ Å will be used subsequently.
The coarse-grained MD simulations demonstrate that damage around the crack is diffused rather than localized, extending into the material beyond the immediate crack tip. The diffusion width $l$ is notably larger than the CG width used to transition from atomic to continuum scales, indicating that the CG width is sufficient to capture local damage and small enough to expose the actual physical spread of damage. Specifically, for fully opened cracks, the damage zone varies in width between approximately 16 and 23 Å, reflecting a non-uniform distribution of damage that suggests its extent may be influenced by structural heterogeneities such as the ring structure, which governs the brittleness and the differences between short-range and medium-range order in glasses [29]. This local variability suggests that the diffusion length of damage is sufficiently small to be influenced by the underlying structural heterogeneity, and can therefore be regarded as a local material property rather than a global parameter.
Finite element scheme.— To show how the MD results can be fed into a continuum scale description, the damage field was then recalculated using a Finite Element Update (FEMU) scheme on 20 slices of the sample to capture the variation along the thickness. During this process, the fracture toughness $g_{c}$ and a constant damage width $l_{c}$ were fitted to minimize the differences between the damage values obtained from coarse-grained molecular scale simulations and those calculated through a phase-field approach [30]. We assumed that the undamaged strain energy density $\psi_{0}$ is known at the atomic scale and minimized the internal energy of the system expressed as:
${\Psi_{{\mathop{\rm int}}}}\left({{\bf{u}},d}\right)=\int_{V}{{\psi_{\epsilon}}\left({{\bf{u}},d}\right)dV}+{g_{c}}\Gamma\left({d,\nabla d}\right),$ (3)
where $\Gamma$ represents the crack surface. In the phase-field approach [31], it is approximated using a crack density function ($\gamma_{\Gamma}$):
$\Gamma=\int\limits_{V}{{\gamma_{\Gamma}}\left({d,\nabla d}\right)dV}=\frac{3}{{8{l_{c}}}}\int_{V}{\left({d+l_{c}^{2}{{\left|{\nabla d}\right|}^{2}}}\right)dV}.$ (4)
Details are provided in the Supplementary Material.
Interestingly, the FEMU produced a damage width $l_{c}$ ranging from 12 to 20 Å, depending on the loading state. This result correlates well with the observed coarse-grained damage width, even though $l_{c}$ is treated as a global parameter in the FEMU. This correlation suggests that, while damage diffusion is locally variable, the global parameter $l_{c}$ effectively captures the average behavior of the damage width. When solving the homogeneous phase-field solution [28], a comparable $\approx 2$ nm regularization length was found based on experimental measurements.
In the literature [21], crack length is often calculated based on an assumed position of the crack tip, either from local density or using a singular Williams series fit. These methods are not sufficiently precise to describe the moving crack. Furthermore, they cannot capture multiple crack fronts. Here the FEMU approach provides a first method to determine the crack length which overcomes these drawbacks.
Free surface energy.—The free surface energy was determined using a smaller cubic sample with a side length of 100 Å. We slice the sample in half by removing atomic bonding along a plane, thus replacing the periodic boundary condition in the direction normal to the plane with a free surface. The mechanical equilibrium of this dissected sample was then reached through potential energy minimization. The free surface energy, $2\gamma$, was calculated by dividing the energy difference by the area of the newly exposed surface. Notably, $2\gamma$ was found to be independent of the sample size once it surpassed the 50 Å box length.
We determined $2\gamma=2.8\pm 0.2$ J/m2. This value is somewhat larger than room temperature experimental measurements including humidity ($\approx 0.5$ J/m2) [11, 32, 33] but in agreement with calculated temperature variations [34] and correlates well with simulations using more sophisticated potentials [21].
We observed that after cutting, equilibrium is reached through atomic displacements localized within a few interatomic distances from the surface. This finding aligns well with experimental observations of surface relaxation in crystals, where an exponential decay over the first few atomic layers has been reported [35]. Detailed information regarding the nature of the energy changes is provided in the Supplementary Material.
[…]
For an elastic deformation, we expect the potential energy $\psi_{\rm{pot}}$ to equal the elastic energy $\psi_{\rm{\epsilon}}$ defined by Eq. 1. Structural evolution however, which is typically not reached through elastic deformation, will be reflected in a variation in $\psi_{\rm{pot}}$. We therefore introduce the difference denoted
$\psi_{\rm{FSE}}=\psi_{\rm{pot}}-\psi_{\rm{\epsilon}}.$ (5)
Fig. 2a shows the distribution of $\psi_{\rm{FSE}}$ along the normal to the surface for the dissected sample. When the coarse-graining is omitted and the energy difference is instead calculated by averaging over 1 Å slices, a pronounced localized peak emerges, as expected for simple surface relaxation. After coarse-graining, the profile width of $\psi_{\rm{SFE}}$ is observed to be 8 Å, reflecting the convolution of the highly localized peak with the width $w$.
A coarse-grained profile for the cracked samples, taken along the line indicated in (b), is also shown. The two coarse-grained $\psi_{\rm{FSE}}$ distributions are similar. This observation suggests that the $\psi_{\rm{FSE}}$ calculated in the fractured sample is the energy change due to the formation of the free surface. It is primarily the result of the surface relaxation due to the loss of atomic connectivity and differs from the damage process shown previously that penetrates significantly deeper into the sample. This is why we refer to this quantity as the free surface energy (FSE) density. Fig. 2b depicts the free surface energy during crack propagation.
Since the energy localization is consistent across different sample conditions, it appears that the free surface energy $2\gamma=\int_{-\infty}^{+\infty}\psi_{\rm{FSE}}\ dy$ should also be an effective measure for assessing the area of newly formed cracks. Fig. 2c displays crack length calculated from the global sample free surface energy $\Psi_{\rm{FSE}}$ divided by $2\gamma$. It is compared with the crack length determined by the FEMU based on the damage approach eq. (4). The two measures are very consistent although the FSE length is smaller than the FEMU length. This agreement between the two methods validates the use of $\psi_{\rm{FSE}}$ as a reliable metric for quantifying crack propagation, providing an alternative to more traditional methods and offering precise measurements that align well with continuum-based models.
Furthermore, analysis of the local structure (see Supplementary Material) revealed that diffuse damage is primarily associated with changes in the ring structure, whereas variations in free surface energy correlate with changes in the coordination number of silicon atoms.
Global energy equilibrium.—As noted in the introduction, experimental measurements have shown that the free surface energy resulting from surface relaxation accounts for only approximately 20% of the energy required to open a crack. This section casts the free surface energy and the damage energy we have calculated at the atomic scale into a global energy balance to identify the origin of this discrepancy.
Fig. 3a illustrates the various energy contributions as a function of loading. The external work, $W_{\rm{ext}}$, is partitioned into several components: the elastic strain energy, $\Psi_{\rm{\epsilon}}$, and the free surface energy (FSE), $\Psi_{\rm{FSE}}$, whose sum is represented as $\Psi_{\rm{pot}}$. Additionally, there is a small non-linear contribution, $\Psi_{\rm{etc}}$, and a significant unquantified remainder. The non-linear energy contribution is determined by isolating the portion of the elastic energy density that exceeds the local potential energy, which ideally remains positive under the assumption of linear elasticity. This excess energy contribution, is then subtracted from the total elastic energy. However, it is noteworthy that this contribution (represented by the dashed line) is minor. This observation suggests that energy dissipation due to shear plasticity, densification or ring folding is indeed negligible. Under tensile loading followed by fracture, these mechanisms contribute only marginally to the overall energy dissipation within the material.
The difference between the external work and the elastic strain energy $W_{\rm{ext}}-\Psi_{\epsilon}$ can be interpreted as the total dissipation caused by fracture, denoted as $\Psi_{\rm{d}}$. This dissipated energy is shown in Fig. 3b (solid red curve) and compared to the dissipation derived from the FEMU fit through eq. 3. Although the two methods are very different (the FEMU does not incorporate explicit knowledge of the external work), they provide very consistent results, highlighting the robustness of the determination of the total energy dissipation. Also shown in the same graph is the FSE, which clearly constitutes only a portion of the total dissipation.
Finally, Fig. 3c presents the fracture toughness ($g_{c}$) as a function of loading, calculated using three distinct methods: the J integral based on the coarse-grained stress and displacement fields (red filled circles – detailed calculations provided in the Supplementary Materials), the FEMU phase-field fit (open circles) and the energy balance approach ($\Psi_{d}/\Gamma_{\rm{SFE}}$) obtained from the coarse-grained atomic scale simulations. The initial linear increase follows from the crack initiation process. Subsequently, the fracture surface energy remains consistent across the three calculation methods. The resulting value provides a ratio of $2\gamma/g_{c}\approx 23$%: this value found at the atomic scale is consistent with experimental findings [11].
Figure 3: Energy balance of the cracked sample as a function the loading parameter  [^3] We note that $K_{I}$ loses its precise physical significance as the tip advances and must be considered as a mere loading parameter.. (a) Various energy contributions normalized by the thickness ($L_{z}$): external work ($W_{\rm{ext}}$), potential energy ($\Psi_{\rm{pot}}$), elastic strain energy ($\Psi_{\rm{\epsilon}}$), surface free energy ($\Psi_{\rm{SFE}}$), and other unquantified contributions (plasticity, folding, non-linear elasticity). (b) Dissipation calculated from the energy balance ($\Psi_{d}$) and crack surface energy from phase-field calculations. (c) Fracture toughness calculated using various methods.
Conclusions.—This Letter sheds new light on the origin of fracture toughness in amorphous silica, an archetype of brittle materials. We investigated the underlying mechanisms that explain the difference between free surface energy and fracture toughness in amorphous silica at the crossover between the atomic and the continuum scales. Using a combination of molecular dynamics and phase-field modeling, we quantified the different contributions to the energy dissipation that occurs during fracture. From a comprehensive energy balance analysis, we were able to clearly isolate the free surface energy term and identify an additional contribution resulting from damage, not plastic deformation. This damage contribution exceeds the free surface energy by a factor of about 4, in agreement with the experimental measurements.
These findings highlight that traditional notions of fracture in brittle materials may overlook the subtle nature of mechanical dissipation in fracture of some amorphous materials. In particular they emphasize the dominant contribution of diffused damage in silicate glasses, aligning with experimental observations. These results open novel perspectives beyond silicate glasses. For example, it would be interesting to evaluate if these concepts are relevant for some metallic glasses [36], which are known to span the transition from brittle to ductile fracture. How they couple to dynamic fracture is also a question worth investigating. The phase-field description is uniquely qualified to identify damage in the presence of plasticity. Therefore, it offers a more nuanced perspective for future materials design, particularly in optimizing the mechanical properties of brittle materials like amorphous silicates where, e.g., the addition of network modifiers may increase the contribution of plasticity.
This research was funded, in part, by French Research National Agency program GaLAaD (ANR-20-CE08-0002).